A flowerpot falls off a windowsill and falls past
Answer #1

Using the kinematic equation to find the speed of the flowerpot when it reaches the window top is

y_2 -y_1 = v_1\cdot t + \frac{1}{2} g t^2

2 = v_1(0.4) + \frac{1}{2} (9.8) (0.4)^2

v_1= 3.04m/s

This is the final velocity of the reach the top of the window from the sill. The time taken for a free fall to reach this speed from the sill is

3.04 - 0 = (9.8) t

t = 0.31s

Now applying the kinematic equation to find the distance of travel from the top of the window to the windowsill is

\Delta y = 0\cdot t + \frac{1}{2} gt^2

\Delta y = \frac{1}{2} (9.8)(0.31)^2

\Delta y = 0.472m